OpenStudy (anonymous):

Calculate the force exerted by 3lb mass in terms of the following. (A) Lb force(english unit)

OpenStudy (anonymous):

\[\Huge\bf \color{yellow}{Welcome~\to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~\to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~\to~\color{purple}{Open}}\color{blue}{Study!!!!}\]

OpenStudy (anonymous):

This question belongs in the physics section, if you go the tab next to openstudy in the upper left and hover over chemistry and click on it, you can search for mire subjects and find physics

OpenStudy (anonymous):

Since you are new, I will help you anyway. The formula for a foot lb is defined as "The pound-force is equal to the gravitational force exerted on a mass of one avoirdupois pound on the surface of Earth"

OpenStudy (anonymous):

In other words \[\Large {\text 1lbf (pound~force)=1 lbm (pound) \times g_n (gravitational force)}\]

OpenStudy (anonymous):

\[\Large \text{1lbf=1lbm x 9.81} \frac{ m }{ s^2 } \]

OpenStudy (anonymous):

\[\Large \text{1lbf=1lbm x 32.17} \frac{ ft }{ s^2 }\] SO just plug in the 3 for lbm and solve for either way you need to express your answer in (meters or feet)